OpenStudy (lexy148):

how to solve this 2 variable linear equation: 3x+2y=12

jhonyy9 (jhonyy9):

do you know the standard form y = mx +c ?

jhonyy9 (jhonyy9):

like a first step rewrite this equation 2y = 12 -3x divide both sides by 2 y = 12/2 -3x/2 y = 6 -3x/2 in the standard form y = -3x/2 +6 so now just give value to x and calculi the y value in this case for example for x=0 y = -3*0/2 +6 y = 6 hope helped you

OpenStudy (mrnood):

it is not possible to 'solve' a two variable system with one equation what do you actually want to do?

OpenStudy (mrnood):

@jhonyy9 has simply re-arranged the equation - but has then taken a 'random' example of x=0, for no particular reason - this is not a solution

OpenStudy (mrnood):

@lexy148 have you written the WHOLE question correctly? (there is no solution the way you posted it)

OpenStudy (mrnood):

 ALL the points on the above line satisfy the given expression